OpenStudy (anonymous):

Calculate the following composition function using the functions f and g defined below: f:N→N, f(x) = x + 5 and g:N→N, g(x) = 5x – 3 (f∘g)(x)

OpenStudy (anonymous):

How do I calculate the value of these function without an x value?

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

f∘g (x) = f ( g(x) ) treat 'g(x)' as the new x . plug it into f f( x) = x + 5 f ( g(x) ) =g(x)+5 replace g(x) by 5x-3 (5x-3) + 5 simplify 5x + 2

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

@freckles is this explanation correct?

OpenStudy (anonymous):

@jim_thompson5910 is this explanation correct?

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

jayzdd has the right explanation

OpenStudy (anonymous):

The online quiz I took said 5x+22; but, I think it was a mistake.

jimthompson5910 (jim_thompson5910):

there has to be a typo

OpenStudy (anonymous):

I'm also trying to figure out where my algebra went wrong on this problem: f(x) = 2x, g(x) = (x-1)/2 I did 2(x-1)/2)) Then 2x-2/2 Finally x-1 The answer this quiz gives me is x-1/2

jimthompson5910 (jim_thompson5910):

They give you `f(x) = 2x, g(x) = (x-1)/2` what exactly are you trying to find? or compute?

OpenStudy (anonymous):

The following functions map ℝ to ℝ. Give an equation describing the composition function g∘f: f(x) = 2x, g(x) = (x-1)/2

jimthompson5910 (jim_thompson5910):